Weight of jack is 3kg more than that of Jill. The weight of both is 16kg less than 213kg.Find the weight of each.
Answer by Rajeev(1)

Let Jill's weight be x kg.
Then, Jack's weight will be x+3 kg.
Thus, x+(x+3)=213-16
or 2x+3 = 197
or 2x = 197-3
i.e. x = 194/2 = 97
Thus, Jill's weight = x = 97 kg
and Jack's weight = x+3 = 97+3 = 100 kg